# HCVP: Leveraging Hierarchical Contrastive Visual Prompt for Domain Generalization

### v1

@@ Abstract
Domain Generalization (DG) endeavors to create machine learning models that excel in unseen scenarios by learning invariant features. In DG, the prevalent practice of constraining models to a fixed structure or uniform parameterization to encapsulate invariant features can inadvertently blend specific aspects. Such an approach struggles with nuanced differentiation of inter-domain variations and may exhibit bias towards certain domains, hindering the precise learning of domain-invariant features. Recognizing this, we introduce a novel method designed to supplement the model with domain-level and task-specific characteristics. This approach aims to guide the model in more effectively separating invariant features from specific characteristics, thereby boosting the generalization. Building on the emerging trend of visual prompts in the DG paradigm, our work introduces the novel Hierarchical Contrastive Visual Prompt (HCVP) methodology. This represents a significant advancement in the field, setting itself apart with a unique generative approach to prompts, alongside an explicit model structure and specialized loss functions. Differing from traditional visual prompts that are often shared across entire datasets, HCVP utilizes a hierarchical prompt generation network enhanced by prompt contrastive learning. These generative prompts are instance-dependent, catering to the unique characteristics inherent to different domains and tasks. Additionally, we devise a prompt modulation network that serves as a bridge, effectively incorporating the generated visual prompts into the vision transformer backbone. Experiments conducted on five DG datasets demonstrate the effectiveness of HCVP, outperforming both established DG algorithms and adaptation protocols.
Index Terms:
Domain Generalization, Visual Prompt, Contrastive Learning
@@ I Introduction
Despite the remarkable successes of machine learning models, particularly deep neural networks (DNNs), in various areas, they often exhibit unexpected failures when the test data distribution deviates from the training data [1, 2, 3]. For instance, adversarial inputs can lead to misclassifications [4], performance on digit recognition may deteriorate under rotations [5], and DNNs trained to recognize pneumonia may fail when applied to scans from new hospitals [6]. Addressing this, domain generalization (DG) seeks to create robust models that can effectively adapt and perform on unseen, yet related, domains by leveraging invariant features learned across multiple source domains [1, 2].
@@
Domain generalization has traditionally relied on techniques such as data manipulation and representation learning [1, 2], giving rise to classic DG algorithms [7, 8, 9, 10, 11, 12]. Despite these efforts, several large-scale empirical studies [13, 14] have highlighted limited performance improvements in current DG algorithms. Alongside these traditional methods, recent advancements in visual foundation models [15], such as vision transformer (ViT) backbone [16], have been leveraged for DG tasks [17, 18]. Their ability to capture a broad spectrum of diverse features has opened up new avenues for effective DG. Several tuning protocols have been proposed to adapt these foundation models to downstream tasks [19, 20, 21, 22, 19, 23].
@@
The fundamental objective in DG lies in effectively discerning and leveraging domain-invariant features from specific characteristics. Nevertheless, the challenge lies in accomplishing this separation within a unified modeling framework without extra, tailored information. The inherent complexity of handling diverse and often subtle differences between domains necessitates additional context beyond what a single model can provide. This lack of specificity leads to a DG approach that unintentionally blends domain-specific aspects, making it challenging to differentiate nuanced inter-domain variations [6]. For example, the DG dataset depicted in Figure 1, encompasses the classification tasks like classifying elephants across diverse domains such as the cartoon, photo, and sketch. Shape attributes like large ears and texture attributes such as thick grey skin are crucial for prediction. However, the significance of the thick grey skin might diminish across domains. Without specific adjustments, a conventional model might struggle to discern these features, leading to the blending of domain-specific attributes in predictions on unseen domains. This complexity highlights the need for a more nuanced approach to handle the unique challenges.
@@
In this work, we introduce a novel method that supplements raw images with domain-level and task-specific characteristics, aiming to guide the model in better learning invariant features from specific aspects. Such insight paves the way for our exploration of visual prompts, which have served as contextual anchors, enriching models with a tailored understanding within specific contexts [24, 25, 26]. Building on the emerging trend of visual prompts in the DG paradigm [27, 28], we introduce the novel Hierarchical Contrastive Visual Prompt (HCVP) methodology, representing a significant advancement in this field. Differing from traditional visual prompts that are often shared across the entire dataset, HCVP employs a two-tier hierarchical prompt generation network along with a prompt contrastive learning strategy [29, 30, 31]. Such explicit model structure and specialized loss functions guide the visual prompt learning process to capture both domain-level and task-specific details. Moreover, we devise a prompt modulation network that acts as a bridge, effectively incorporating our generated visual prompts into the ViT backbone. In order to enhance visual features with the discriminative power between inter-classes, we introduce class-conditioned contrastive invariance. The experiments, conducted on several DG datasets [13, 32], demonstrate the effectiveness of our approach.
@@ II-A Domain Generalization
Domain generalization traditionally relied on data manipulation and representation learning [1]. Techniques such as data augmentation [33] and generation [34] were utilized to diversify training data. Representation learning [35, 36, 37, 38, 39, 40, 41, 42] used strategies ranging from feature alignment [43], domain adversarial learning [44, 7], to causality-inspired methods [8, 9, 10]. Recent studies suggest limited improvement in DG methods over empirical risk minimization (ERM) [13, 14], indicating a need for innovation. While visual foundation models are employed in modern DG techniques [17, 15], they often necessitate extensive model selection [20, 21]. Various protocols have been proposed for adaptation, such as linear probing [45, 22], LP-FT (linear probing and then full fine-tuning) [19], and gradual unfreeze (fine-tuning last $k$ layers) [23, 46]. However, conventional DG methods typically rely on a universal parameter set and my blend domain-specific features, and thus hinder generalization. In contrast, our approach focuses on enriching the model with domain-level and task-specific characteristics to enhance separation of invariant features.
@@ II-B Visual Prompt Tuning
Prompt tuning acts as a powerful tool that furnishes transformer-based large models with the necessary semantic context for particular tasks [47, 48]. Such a principle has successfully transitioned to the visual realm, culminating in visual prompt tuning (VPT) [26]. Traditional visual prompt methods [26, 49] typically implement visual prompts using parameter placeholders, where a single set of prompts is shared across various tasks. While effective for specific downstream tasks [50, 51], they are not specifically tailored for OoD scenarios, potentially hindering their ability to bridge distribution shifts within the DG context. Interestingly, the potential benefits of more customized prompts have been observed in other areas, such as the use of class-specific prompts for certain fine-grained categories [25]. This observation highlights the potential of adapting visual prompts to embody domain-level and task-specific details within DG. Recent works, namely DoPrompt [27] and CSVPT [28], have begun to explore the application of visual prompting within DG paradigms. However, our HCVP presents distinct advancements in this area. Unlike DoPrompt [27], which uses prompts shared across each domain and leads to challenges in prompt selection or fusion, HCVP utilizes generative prompts that encapsulate more nuanced characteristics. Besides, HCVP eliminates the need for the extra adapters required in DoPrompt. Moreover, contrasting with CSVPT [28], which also adopts generative prompting but lacks additional regularization mechanisms, HCVP introduces an explicit model structure and specialized loss functions. These are meticulously designed to guide the visual prompt learning process, ensuring a more targeted and effective integration of domain-specific and task-specific information within the DG framework.
@@ III-A Notations
To establish a clear understanding of our approach, we define several critical variables that are central to our methodology: $Y$ (the class label we aim to predict), $D$ (the domain index), $Z$ (the learned invariant features crucial for domain generalization), $X$ (the raw input features), and $P$ (the learned prompts representing specific characteristics extracted from $X$). Following [1, 2], we denote the input feature space as $\mathcal{X}$ and the target label space as $\mathcal{Y}$. A domain comprises data from a joint distribution $P_{XY}$ in the space $\mathcal{X}\times\mathcal{Y}$. For a given domain $P_{XY}$, $P_{X}$, $P_{Y|X}$, and $P_{X|Y}$ represent the marginal, posterior, and class-conditional distributions, respectively. We have access to $M$ source domains $\mathcal{S}_{train}=\{\mathcal{S}^{i}|i=1,\cdots,M\}$, each $\mathcal{S}^{i}=\{(\bm{x}_{j}^{i},y_{j}^{i})\}_{j=1}^{n_{i}}$, where $\bm{x}$ and $y$ denote the vectors of input features and class labels. The joint distributions between any two domains are not equal, i.e., $P_{XY}^{i}\neq P_{XY}^{j},1\leq i\neq j\leq M$. The objective in DG is to learn a function $f:\mathcal{X}\rightarrow\mathcal{Y}$ using the $M$ training domains to minimize prediction error on a new test domain $\mathcal{T}=\{\bm{x}^{\mathcal{T}},y^{\mathcal{T}}\}$. In our work, this function $f$ is defined as $f(\bm{x},\bm{p})$, where $\bm{p}$ represents generative visual prompts that are integrated into the model to enhance its learning capability:
[…]
@@ III-B Role of Generative Visual Prompts in Mutual Information
Mutual information is a measure from information theory that quantifies the amount of information obtained about one random variable through observing another random variable [52, 53]. In the context of DG, we consider two key mutual information terms [54]:
@@ IV-A Hierarchical Prompt Generation Network
As discussed in Sec. II-B, although traditional visual prompt methods [26] are effective for specific downstream tasks, they are not specifically tailored for OoD scenarios. While merely replicating visual prompts for individual domain can allow prompts to mirror domain-level information, prompt selection or fusion becomes challenging in DG settings where the target domain is unknown or inaccessible [27]. To tackle these challenges, our Hierarchical Prompt Generation Network (HPGN) offers a sophisticated solution. While CSVPT [28] also utilizes generative visual prompts, HPGN, goes a step further by integrating prompt contrastive learning. In our work, the model structure and loss functions are specifically designed to generate visual prompts that adeptly encapsulate domain-level and task-specific information.
@@ IV-B Prompt Modulation Network
Within our architecture, the Prompt Modulation Network (PMN) serves as a conduit to effectively integrate the generated visual prompts into the ViT backbone. Unlike traditional visual prompt methods [26] that rely on fixed prompt placeholders, our use of dynamically generated prompts necessitates this specially designed integration mechanism. Specifically, the PMN receives the two-tier generated visual prompts as input and transforming them through a MLP. Importantly, the MLP is configured with a number of layers that is identical to the layer count in the ViT. This transformation process is conducted during training, where the transformed prompts are systematically integrated into each ViT layer $L_{i}$, to facilitate a nuanced adaptation of the visual content. The class token vector and patch embeddings after $i$-th ViT layer are represented by $\mathbf{x}_{i}$ and $\mathbf{E}_{i}$, and the integration is formalized as:
@@
This modulation mechanism, drawing on principles from batch normalization [56] and attention mechanisms [57, 58], allows the PMN to adapt the ViT’s internal representations, effectively utilizing the generated visual prompts. This enhances the model’s ability to accommodate both domain-level and task-specific details.
@@ IV-E End-to-End Training Strategy
Our training strategy combines multiple loss components: prompt contrastive learning loss $\mathcal{L}_{\text{PCL}}$, class-conditioned contrastive invariance loss $\mathcal{L}_{\text{CCI}}$, and classification loss $\mathcal{L}_{\text{cls}}$. The classification loss is defined using cross-entropy between the output from the classification head applied to $\mathbf{x}_{N}$ and the ground-truth label $y$, aiming to increase the mutual information term $I(Z,Y)$:

### v2

@@ Abstract
Domain Generalization (DG) endeavors to create machine learning models that excel in unseen scenarios by learning invariant features. In DG, the prevalent practice of constraining models to a fixed structure or uniform parameterization to encapsulate invariant features can inadvertently blend specific aspects. Such an approach struggles with nuanced differentiation of inter-domain variations and may exhibit bias towards certain domains, hindering the precise learning of domain-invariant features. Recognizing this, we introduce a novel method designed to supplement the model with domain-level and task-specific characteristics. This approach aims to guide the model in more effectively separating invariant features from specific characteristics, thereby boosting the generalization. Building on the emerging trend of visual prompts in the DG paradigm, our work introduces the novel Hierarchical Contrastive Visual Prompt (HCVP) methodology. This represents a significant advancement in the field, setting itself apart with a unique generative approach to prompts, alongside an explicit model structure and specialized loss functions. Differing from traditional visual prompts that are often shared across entire datasets, HCVP utilizes a hierarchical prompt generation network enhanced by prompt contrastive learning. These generative prompts are instance-dependent, catering to the unique characteristics inherent to different domains and tasks. Additionally, we devise a prompt modulation network that serves as a bridge, effectively incorporating the generated visual prompts into the vision transformer backbone. Experiments conducted on five DG datasets demonstrate the effectiveness of HCVP, outperforming both established DG algorithms and adaptation protocols. The code is available at: https://github.com/jameszhou-gl/TMM-HCVP.
@@ I Introduction
Despite the remarkable successes of machine learning models, particularly deep neural networks (DNNs), in various areas, they often exhibit unexpected failures when the test data distribution deviates from the training data [wang2022generalizing, zhou2022domain, zhou2023emerging]. For instance, adversarial inputs can lead to misclassifications [goodfellow2014explaining], performance on digit recognition may deteriorate under rotations [piratla2020efficient], and DNNs trained to recognize pneumonia may fail when applied to scans from new hospitals [geirhos2020shortcut]. Addressing this, domain generalization (DG) seeks to create robust models that can effectively adapt and perform on unseen, yet related, domains by leveraging invariant features learned across multiple source domains [wang2022generalizing, zhou2022domain].
@@
Domain generalization has traditionally relied on techniques such as data manipulation and representation learning [wang2022generalizing, zhou2022domain], giving rise to classic DG algorithms [ganin2016domain, rojas2018invariant, sun2021recovering, christiansen2021causal, li2018deep, wang2023causal]. Despite these efforts, several large-scale empirical studies [gulrajani2020search, wiles2022a] have highlighted limited performance improvements in current DG algorithms. Alongside these traditional methods, recent advancements in visual foundation models [bommasani2021opportunities], such as vision transformer (ViT) backbone [dosovitskiy2021an], have been leveraged for DG tasks [zhang2022delving, chen2023explore]. Their ability to capture a broad spectrum of diverse features has opened up new avenues for effective DG. Several tuning protocols have been proposed to adapt these foundation models to downstream tasks [kumar2022fine, li2022sparse, rame2023model, du2021fewshot, kumar2022fine, lee2022surgical].
@@
The fundamental objective in DG lies in effectively discerning and leveraging domain-invariant features from specific characteristics. Nevertheless, the challenge lies in accomplishing this separation within a unified modeling framework without extra, tailored information. The inherent complexity of handling diverse and often subtle differences between domains necessitates additional context beyond what a single model can provide. This lack of specificity leads to a DG approach that unintentionally blends domain-specific aspects, making it challenging to differentiate nuanced inter-domain variations [geirhos2020shortcut]. For example, the DG dataset depicted in Figure 1, encompasses the classification tasks like classifying elephants across diverse domains such as the cartoon, photo, and sketch. Shape attributes like large ears and texture attributes such as thick grey skin are crucial for prediction. However, the significance of the thick grey skin might diminish across domains. Without specific adjustments, a conventional model might struggle to discern these features, leading to the blending of domain-specific attributes in predictions on unseen domains. This complexity highlights the need for a more nuanced approach to handle the unique challenges.
@@
In this work, we introduce a novel method that supplements raw images with domain-level and task-specific characteristics, aiming to guide the model in better learning invariant features from specific aspects. Such insight paves the way for our exploration of visual prompts, which have served as contextual anchors, enriching models with a tailored understanding within specific contexts [radford2021learning, zhou2022learning, jia2022visual]. Building on the emerging trend of visual prompts in the DG paradigm [zheng2022prompt, li2022learning], we introduce the novel Hierarchical Contrastive Visual Prompt (HCVP) methodology, representing a significant advancement in this field. Differing from traditional visual prompts that are often shared across the entire dataset, HCVP employs a two-tier hierarchical prompt generation network along with a prompt contrastive learning strategy [chen2020simple, zhou2023contrastive, li2022contrastive]. Such explicit model structure and specialized loss functions guide the visual prompt learning process to capture both domain-level and task-specific details. Moreover, we devise a prompt modulation network that acts as a bridge, effectively incorporating our generated visual prompts into the ViT backbone. In order to enhance visual features with the discriminative power between inter-classes, we introduce class-conditioned contrastive invariance. The experiments, conducted on several DG datasets [gulrajani2020search, ye2022ood], demonstrate the effectiveness of our approach.
@@ II-A Domain Generalization
Domain generalization traditionally relied on data manipulation and representation learning [wang2022generalizing]. Techniques such as data augmentation [yue2019domain] and generation [qiao2020learning] were utilized to diversify training data. Representation learning [jo2023poem, cha2021swad, cha2022domain, liu2021domain, jin2021style, ma2019deep, luo2023taking, niu2023knowledge, wang2021learning, gokhale2023improving] used strategies ranging from feature alignment [li2018domain], domain adversarial learning [ganin2015unsupervised, ganin2016domain], to causality-inspired methods [rojas2018invariant, sun2021recovering, christiansen2021causal]. Recent studies suggest limited improvement in DG methods over empirical risk minimization (ERM) [gulrajani2020search, wiles2022a], indicating a need for innovation. While visual foundation models are employed in modern DG techniques [zhang2022delving, bommasani2021opportunities, han2023well], they often necessitate extensive model selection [li2022sparse, rame2023model]. Various protocols have been proposed for adaptation, such as linear probing [Wu2020Understanding, du2021fewshot], LP-FT (linear probing and then full fine-tuning) [kumar2022fine], and gradual unfreeze (fine-tuning last $k$ layers) [lee2022surgical, he2022masked]. However, conventional DG methods typically rely on a universal parameter set and my blend domain-specific features, and thus hinder generalization. In contrast, our approach focuses on enriching the model with domain-level and task-specific characteristics to enhance separation of invariant features.
@@ II-B Visual Prompt Tuning
Prompt tuning acts as a powerful tool that furnishes transformer-based large models with the necessary semantic context for particular tasks [Li2021PrefixTuningOC, Liu2021PTuningVP]. Such a principle has successfully transitioned to the visual realm, culminating in visual prompt tuning (VPT) [jia2022visual]. Traditional visual prompt methods [jia2022visual, chen2023understanding] typically implement visual prompts using parameter placeholders, where a single set of prompts is shared across various tasks. While effective for specific downstream tasks [Sohn2022VisualPT, Das2023LearningEP], they are not specifically tailored for OoD scenarios, potentially hindering their ability to bridge distribution shifts within the DG context. Interestingly, the potential benefits of more customized prompts have been observed in other areas, such as the use of class-specific prompts for certain fine-grained categories [zhou2022learning]. This observation highlights the potential of adapting visual prompts to embody domain-level and task-specific details within DG. Recent works, namely DoPrompt [zheng2022prompt] and CSVPT [li2022learning], have begun to explore the application of visual prompting within DG paradigms. However, our HCVP presents distinct advancements in this area. Unlike DoPrompt [zheng2022prompt], which uses prompts shared across each domain and leads to challenges in prompt selection or fusion, HCVP utilizes generative prompts that encapsulate more nuanced characteristics. Besides, HCVP eliminates the need for the extra adapters required in DoPrompt. Moreover, contrasting with CSVPT [li2022learning], which also adopts generative prompting but lacks additional regularization mechanisms, HCVP introduces an explicit model structure and specialized loss functions. These are meticulously designed to guide the visual prompt learning process, ensuring a more targeted and effective integration of domain-specific and task-specific information within the DG framework. In domain adaptation (DA) area, DAPrompt [ge2023domain] utilizes textual prompt information which requires manual design while our HCVP introduces an innovative approach by generating visual prompts in an end-to-end manner.
@@ III-A Notations
To establish a clear understanding of our approach, we define several critical variables that are central to our methodology: $Y$ (the class label we aim to predict), $D$ (the domain index), $Z$ (the learned invariant features crucial for domain generalization), $X$ (the raw input features), and $P$ (the learned prompts representing specific characteristics extracted from $X$). Following [wang2022generalizing, zhou2022domain], we denote the input feature space as $\mathcal{X}$ and the target label space as $\mathcal{Y}$. A domain comprises data from a joint distribution $P_{XY}$ in the space $\mathcal{X}\times\mathcal{Y}$. For a given domain $P_{XY}$, $P_{X}$, $P_{Y|X}$, and $P_{X|Y}$ represent the marginal, posterior, and class-conditional distributions, respectively. We have access to $M$ source domains $\mathcal{S}_{train}=\{\mathcal{S}^{i}|i=1,\cdots,M\}$, each $\mathcal{S}^{i}=\{(\bm{x}_{j}^{i},y_{j}^{i})\}_{j=1}^{n_{i}}$, where $\bm{x}$ and $y$ denote the vectors of input features and class labels. The joint distributions between any two domains are not equal, i.e., $P_{XY}^{i}\neq P_{XY}^{j},1\leq i\neq j\leq M$. The objective in DG is to learn a function $f:\mathcal{X}\rightarrow\mathcal{Y}$ using the $M$ training domains to minimize prediction error on a new test domain $\mathcal{T}=\{\bm{x}^{\mathcal{T}},y^{\mathcal{T}}\}$. In our work, this function $f$ is defined as $f(\bm{x},\bm{p})$, where $\bm{p}$ represents generative visual prompts that are integrated into the model to enhance its learning capability:
$\underset{f}{\text{min}}\mathbb{E}_{(\bm{x},y)\in\mathcal{T}}[l(f(\bm{x},\bm{p}),y)]$ (1)
@@ III-B Role of Generative Visual Prompts in Mutual Information
Mutual information is a measure from information theory that quantifies the amount of information obtained about one random variable through observing another random variable [tishby2000information, tishby2015deep]. In the context of DG, we consider two key mutual information terms [li2022invariant]:
@@ IV-A Hierarchical Prompt Generation Network
As discussed in Sec. II-B, although traditional visual prompt methods [jia2022visual] are effective for specific downstream tasks, they are not specifically tailored for OoD scenarios. While merely replicating visual prompts for individual domain can allow prompts to mirror domain-level information, prompt selection or fusion becomes challenging in DG settings where the target domain is unknown or inaccessible [zheng2022prompt]. To tackle these challenges, our Hierarchical Prompt Generation Network (HPGN) offers a sophisticated solution. While CSVPT [li2022learning] also utilizes generative visual prompts, HPGN, goes a step further by integrating prompt contrastive learning. In our work, the model structure and loss functions are specifically designed to generate visual prompts that adeptly encapsulate domain-level and task-specific information.
@@ IV-B Prompt Modulation Network
Within our architecture, the Prompt Modulation Network (PMN) serves as a conduit to effectively integrate the generated visual prompts into the ViT backbone. Unlike traditional visual prompt methods [jia2022visual] that rely on fixed prompt placeholders, our use of dynamically generated prompts necessitates this specially designed integration mechanism. Specifically, the PMN receives the two-tier generated visual prompts as input and transforming them through a MLP. Importantly, the MLP is configured with a number of layers that is identical to the layer count in the ViT. This transformation process is conducted during training, where the transformed prompts are systematically integrated into each ViT layer $L_{i}$, to facilitate a nuanced adaptation of the visual content. The class token vector and patch embeddings after $i$-th ViT layer are represented by $\mathbf{x}_{i}$ and $\mathbf{E}_{i}$, and the integration is formalized as:
@@
This modulation mechanism, drawing on principles from batch normalization [ioffe2015batch] and attention mechanisms [bahdanau2015neural, vaswani2017attention], allows the PMN to adapt the ViT’s internal representations, effectively utilizing the generated visual prompts. This enhances the model’s ability to accommodate both domain-level and task-specific details.
@@ IV-E End-to-End Training Strategy
Our training strategy combines multiple loss components: prompt contrastive learning loss $\mathcal{L}_{\text{PCL}}$, class-conditioned contrastive invariance loss $\mathcal{L}_{\text{CCI}}$, and classification loss $\mathcal{L}_{\text{cls}}$, as shown in Algorithm 1. The classification loss is defined using cross-entropy between the output from the classification head applied to $\mathbf{x}_{N}$ and the ground-truth label $y$, aiming to increase the mutual information term $I(Z,Y)$:
@@
Algorithm 1 Hierarchical Contrastive Visual Prompt (HCVP) Algorithm
1:Number of steps $T$, pretrained ResNet encoder, Vision Transformer (ViT) model $f$, classifier $\phi$, and hyper-parameters $\lambda_{PCL}$, $\lambda_{CCI}$.
2:Initialize Hierarchical Prompt Generation Network (HPGN) and Prompt Modulation Network (PMN) parameters.
3:for $t\leftarrow 1$ to $T$ do
4:  Sample a batch of images and labels from $M$ source domains $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{M}\}$:
5:  for $i\leftarrow 1$ to $M$ do
6:   $\{(\bm{x}_{j}^{i},y_{j}^{i})\}_{j=1}^{n_{i}}\leftarrow$ Sample from domain $\mathcal{S}_{i}$
7:  end for
8:  Generate domain and task prompts using HPGN based on Eq. (3) and Eq. (4).
9:  Modulate prompts with PMN and integrate into ViT layers as per Eq. (5) and Eq. (6).
10:  Calculate Prompt Contrastive Learning (PCL) loss $\mathcal{L}_{PCL}$ using Eq. (8) and Eq. (10).
11:  Compute Class-Conditional Contrastive Invariance (CCI) loss $\mathcal{L}_{CCI}$ using Eq. (11).
12:  Compute classification loss $\mathcal{L}_{cls}$ using Eq. (12).
13:  Total loss $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{cls}}+\lambda_{PCL}\mathcal{L}_{\text{PCL}}+\lambda_{CCI}\mathcal{L}_{\text{CCI}}$.
14:  Backward $\mathcal{L}_{\text{total}}$ and update the parameters.
15:end for
16:return Trained HCVP model and results based on the model selection criteria using Training-domain validation as per [gulrajani2020search].